# Chains of model structures arising from modules of finite Gorenstein dimension

### v2

@@ Introduction
Gorenstein projective modules, introduced by M. Auslander and M. Bridger [1], and E. Enochs and O. M. G. Jenda [10], have got deep interests and wide applications in mathematics. A recent milestone is the introduction of projectively coresolved Gorenstein flat (PGF, for short) modules over an arbitrary ring $R$, by J. Šaroch and J. Št́ovíček [28].
@@
What interesting and important are, all the three classes $\mathcal{PGF}$, $\mathcal{GP}$, and $\mathcal{GF}$, are the left parts of cotorsion pairs in $R$-Mod, where $\mathcal{GP}$ (respectively, $\mathcal{GF}$) is the full subcategory of $R$-Mod consisting of the Gorenstein projective (respectively, Gorenstein flat) modules. A. Beligiannis and I. Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod, if $A$ is an Artin algebra. Šaroch and Št́ovíček [28, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$. By a result of A. Iacob [20, Proposition 9] (see also [22, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$. Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [19, Theorem 8.3]. By the Hovey correspondence, all these three kinds of cotorsion pairs induce abelian model structures in $R$-Mod.
@@
is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra, where $\mathcal{GP}_{n}$ is the class of $R$-modules of Gorenstein projective dimension $\leq n$ (see Theorem 3.4). R. EI Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
@@
which are not trivial (i.e., not every module is a trivial object in the model structure), and not projective (i.e., not every module is a fibrant object) in general. (We remind that, in this paper, a Hovey triple is written in the order  $(\mathcal{C},\mathcal{F},\mathcal{W})$, rather than  $(\mathcal{C},\mathcal{W},\mathcal{F})$, where $\mathcal{W}$ is the class of trivial objects.) Since all the hereditary abelian model structures in each chain have the same class of trivial objects, the homotopy categories in the same chain are triangle equivalent to each other. Thus the homotopy categories are the stable categories $\mathcal{PGF}/\mathcal{P}$ and $\mathcal{GP}/\mathcal{P}$.
@@
and one chain of hereditary Hovey triples $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ in $\mathcal{GF}_{n}$. See Corollary 6.1. In this way one gets double chains of complete and hereditary cotorsion pairs and hereditary Hovey triples, parameterized by $m$ and $n$. When $m=0=n$, this is G. Dalezios and I. Emmanouil [8, Theorem 4.3]. All these exact model structures are not trivial, and not projective in general, with the same homotopy category $\mathcal{PGF}/\mathcal{P}$.
@@ 1.2. Cotorsion pairs in exact categories
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$ for a set $\mathcal{S}$. Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([12, Theorem 10]; also [16, Theorem 6.11(b)]: but in [16] “cogenerated” is called “generated”). This result has been generalized to Grothendieck category with enough projective objects in [19, Theorem 2.4].
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is hereditary, if $\mathcal{C}$ is closed under the kernel of deflations, and $\mathcal{F}$ is closed under the cokernel of inflations. For a cotorsion pair $(\mathcal{C},\mathcal{F})$ in an abelian category $\mathcal{A}$ with enough projective objects and injective objects, these two conditions of the heredity are equivalent; and they are equivalent to ${{\rm Ext}}^{2}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$ and $F\in\mathcal{F};$ and also to ${{\rm Ext}}^{i}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$, $F\in\mathcal{F}$, and $i\geq 2$. See [13, Proposition 1.2.10]. This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category. See [27, Lemma 6.17].
@@ 1.3. Model structures
A closed model structure, or simply, a model structure, on a category $\mathcal{M}$, is a triple  $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ of classes of morphisms, where the morphisms in the three classes are respectively called cofibrations, fibrations, and weak equivalences, satisfying Two out of three axiom, Retract axiom, Lifting axiom, and Factorization axiom. The morphisms in $\operatorname{CoFib}\cap\operatorname{Weq}$ (respectively, $\operatorname{Fib}\cap\operatorname{Weq}$) are called trivial cofibrations (respectively, trivial fibrations). For details we refer to [24], or [18]. In a model structure any two classes of $\operatorname{CoFib}$,   $\operatorname{Fib}$,   $\operatorname{Weq}$ uniquely determine the third one ([25]).
@@
M. Hovey [19] has established a one-one correspondence between abelian model structures and the Hovey triples in an abelian category. This has been extended to weakly idempotent complete exact category in J. Gillespie [14, Theorem 3.3] (see also [27, Theorem 6.9]). Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
@@
${\rm Weq}=\{pi\mid i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{C}\cap\mathcal{W},\ p\ \mbox{is a deflation,}\ {\rm Ker}p\in\mathcal{F}\cap\mathcal{W}\}.$
@@ 1.4. Homotopy categories of exact model structures
A morphism $f$ in exact category is admissible if $f=ip$, where $p$ is a deflation and $i$ is an inflation. In this case, ${\rm Ker}f$ and ${\rm Coker}f$ exist, and ${\rm Ker}f={\rm Ker}p$ and ${\rm Coker}f={\rm Coker}i$.
@@
$(1)$   ([27, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
@@
For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([27, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5]. The following nice result can be explicitly founded in Št́ovíček [27, Theorem 6.21]. See also Gillespie [14, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
@@
By Corollary 1.2, one only needs to show that the two triangulations in the same homotopy category also coincide. This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$. See [27, Definition 6.15]. ∎
@@ 2. Abelian model structures arising from $\mathcal{PGF}_{n}$
([8, Proposition 3.1])  Let $R$ be a ring, and $M$ an $R$-module with ${\rm PGFd}M=n$. Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and ${\rm pd}K=n-1$. (normal-(((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{PGF}$-approximation of $M$.
@@
Let $R$ be a ring, $n$ a non-negative integer, and $M$ an $R$-module. Then $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n};$ and the following are equivalent:normal-:::
@@
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$. (normal-(((If $n=0$, then $K=0.)$
@@ 3. Abelian model structures arising from $\mathcal{GP}_{n}$
([5, X, Theorem 2.4(iv)])  Let $A$ be an Artin algebra. Then  $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod. Moreover, $\mathcal{GP}^{\perp}$ is thick in $A$-Mod (normal-(((this does not need the assumption of Artin algebra)normal-))).
@@
Let $R$ be a ring, and $M$ an $R$-module with ${\rm Gpd}M=n$. Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{GP}$ and ${\rm pd}K=n-1$. (normal-(((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{GP}$-approximation of $M$.
@@
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. (normal-(((If $n=0$, then $K=0.)$
@@ 5.1. Induced complete cotorsion pairs in exact categories
$(2)$  Since $\mathcal{X}$ is closed under the kernels of epimorphisms in $\mathcal{A}$, $\mathcal{X}$ is closed under the kernels of deflations in $\mathcal{B}$. Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf. [27, Lemma 6.17]; or Subsection 1.2). ∎
@@ 6.3. A sufficient and necessary condition for $\mathcal{PGF}=\mathcal{GP}$
Let $R$ be a ring. Then the following are equivalent:normal-:::
@@ 7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
The following result is due to R. EI Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
@@
([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent:normal-:::
@@
When $n=0$, the following result is [28, Theorem 4.9].
Theorem 7.2.
@@
where $0\leq m\leq n$ (see Lemma 2.2, Lemma 7.3, Corollary 2.6, and Theorem 7.2). Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GF}_{n}$, one gets the assertions (1) - (4) in the following corollary.

### v3

@@ Introduction
Gorenstein projective modules, introduced by M. Auslander and M. Bridger [1], and E. Enochs and O. M. G. Jenda [10], have got deep interests and wide applications in mathematics. A recent important progress is the introduction of projectively coresolved Gorenstein flat (PGF, for short) modules over an arbitrary ring $R$, by J. Šaroch and J. Št́ovíček [28].
@@
What interesting and important are, all the three classes $\mathcal{PGF}$, $\mathcal{GP}$, and $\mathcal{GF}$, are the left parts of cotorsion pairs in $R$-Mod, where $\mathcal{GP}$ (respectively, $\mathcal{GF}$) is the full subcategory of $R$-Mod consisting of the Gorenstein projective (respectively, Gorenstein flat) modules. A. Beligiannis and I. Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra. Šaroch and Št́ovíček [28, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$. By a result of A. Iacob [20, Proposition 9] (see also [22, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$. Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [19, Theorem 8.3]. By the Hovey correspondence, all these three kinds of cotorsion pairs induce abelian model structures in $R$-Mod.
@@
is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra, where $\mathcal{GP}_{n}$ is the class of $R$-modules of Gorenstein projective dimension $\leq n$ (see Theorem 3.4). R. El Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
@@
which are not trivial (i.e., not every module is a trivial object in the model structure), and not projective (i.e., not every module is a fibrant object) in general. (We remind that, in this paper, a Hovey triple is written in the order  $(\mathcal{C},\mathcal{F},\mathcal{W})$, rather than  $(\mathcal{C},\mathcal{W},\mathcal{F})$, where $\mathcal{W}$ is the class of trivial objects.) Since all the hereditary abelian model structures in each chain have the same class of trivial objects, the homotopy categories in the same chain are triangle equivalent to each other. Thus the homotopy categories are the stable categories $\mathcal{PGF}/\mathcal{P}$ and $\mathcal{GP}/\mathcal{P}$, respectively.
@@
and one chain of hereditary Hovey triples $(\mathcal{PGF}_{m},\ \mathcal{P}_{m}^{\perp}\cap\mathcal{GP}_{n},\ \mathcal{PGF}^{\perp}\cap\mathcal{GP}_{n})$ in $\mathcal{GP}_{n}$. See Corollary 6.1. In this way one gets double chains of complete and hereditary cotorsion pairs and hereditary Hovey triples, parameterized by $m$ and $n$. All these exact model structures are not trivial, and not projective in general, with the same homotopy category $\mathcal{PGF}/\mathcal{P}$.
@@ 1.2. Cotorsion pairs in exact categories
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$, where $\mathcal{S}$ is a set of modules (or, equivalently, cogenerated by a single module). Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([12, Theorem 10]; also [16, Theorem 6.11(b)]: but in [16] “cogenerated” is called “generated”). This result has been generalized to Grothendieck category with enough projective objects in [19, Theorem 2.4].
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is hereditary, if $\mathcal{C}$ is closed under the kernel of deflations, and $\mathcal{F}$ is closed under the cokernel of inflations. For a cotorsion pair $(\mathcal{C},\mathcal{F})$ in an abelian category $\mathcal{A}$ with enough projective objects and injective objects, these two conditions of the heredity are equivalent; and they are equivalent to ${{\rm Ext}}^{2}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$ and $F\in\mathcal{F};$ and also to ${{\rm Ext}}^{i}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$, $F\in\mathcal{F}$, and $i\geq 2$. See [13, Proposition 1.2.10]. This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category. See [29, Lemma 6.17] (also [27]).
@@ 1.3. Model structures
A closed model structure, or simply, a model structure, on a category $\mathcal{M}$, is a triple  $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ of classes of morphisms, where the morphisms in the three classes are respectively called cofibrations, fibrations, and weak equivalences, satisfying Two out of three axiom, Retract axiom, Lifting axiom, and Factorization axiom. The morphisms in ${\rm TCoFib}:=\operatorname{CoFib}\cap\operatorname{Weq}$ (respectively, ${\rm TFib}:=\operatorname{Fib}\cap\operatorname{Weq}$) are called trivial cofibrations (respectively, trivial fibrations). For details we refer to [24], or [18]. In a model structure any two classes of $\operatorname{CoFib}$,   $\operatorname{Fib}$,   $\operatorname{Weq}$ uniquely determine the third one ([25]).
@@
M. Hovey [19] has established a one-one correspondence between abelian model structures and the Hovey triples in an abelian category. This has been extended to weakly idempotent complete exact category in J. Gillespie [14, Theorem 3.3] (see also [29, Theorem 6.9]). Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
@@
${\rm Weq}=\{pi\mid i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{C}\cap\mathcal{W},\ p\ \mbox{is a deflation,}\ {\rm Ker}p\in\mathcal{F}\cap\mathcal{W}\}$
and
${\rm TCoFib}=\{\mbox{inflation with cokernel in}\ \mathcal{C}\cap\mathcal{W}\},\ \ \ \ {\rm TFib}=\{\mbox{deflation with kernel in}\ \mathcal{F}\cap\mathcal{W}\}.$
@@ 1.4. Homotopy categories of exact model structures
A morphism $f$ in exact category is admissible if $f=ip$, where $p$ is a deflation and $i$ is an inflation. In this case, ${\rm Ker}f$ and ${\rm Coker}f$ exist, and ${\rm Ker}f={\rm Ker}p$ and ${\rm Coker}f={\rm Coker}i$. The following description of $\operatorname{Weq}$ in $(2)$ is similar to the multiplicative system determined by a Serre subcategory of an abelian category.
@@
$(1)$   ([29, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
@@
For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([29, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5]. The following nice result can be explicitly founded in Št́ovíček [29, Theorem 6.21]. See also Gillespie [14, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
@@
By Corollary 1.2, one only needs to show that the two triangulations in the same homotopy category also coincide. This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$. See [29, Definition 6.15]. ∎
@@ 2. Abelian model structures arising from $\mathcal{PGF}_{n}$
([8, Proposition 3.1])  Let $R$ be a ring, and $M$ an $R$-module with ${\rm PGFd}M=n$. Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{PGF}$ and ${\rm pd}K=n-1$. ((((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{PGF}$-approximation of $M$.
@@
Let $R$ be a ring, $n$ a non-negative integer, and $M$ an $R$-module. Then $\mathcal{PGF}^{\perp}\cap\mathcal{PGF}_{n}=\mathcal{P}_{n};$ and the following are equivalent::::
@@
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{R}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{PGF}$ and $K\in\mathcal{P}_{n-1}$. ((((If $n=0$, then $K=0.)$
@@ 3. Abelian model structures arising from $\mathcal{GP}_{n}$
([5, X, Theorem 2.4(iv)])  Let $A$ be an Artin algebra. Then  $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $A$-Mod. Moreover, $\mathcal{GP}^{\perp}$ is thick in $A$-Mod ((((this does not need the assumption of Artin algebra)))).
@@
Let $R$ be a ring, and $M$ an $R$-module with ${\rm Gpd}M=n$. Then there is an exact sequence $0\longrightarrow K\longrightarrow G\stackrel{{\scriptstyle\phi}}{{\longrightarrow}}M\longrightarrow 0$ with $G\in\mathcal{GP}$ and ${\rm pd}K=n-1$. ((((If $n=0$ then $K=0.)$ In particular, $\phi$ is a right $\mathcal{GP}$-approximation of $M$.
@@
(2)  There is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ which is again exact after applying $\operatorname{Hom}_{A}(-,X)$ for $X\in\mathcal{P}_{n}^{\perp}$, where $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. ((((If $n=0$, then $K=0.)$
@@ 5.1. Induced complete cotorsion pairs in exact categories
$(2)$  Since $\mathcal{X}$ is closed under the kernels of epimorphisms in $\mathcal{A}$, $\mathcal{X}$ is closed under the kernels of deflations in $\mathcal{B}$. Since $(\mathcal{X},\ \mathcal{Y}\cap\mathcal{B})$ is a complete cotorsion pair in $\mathcal{B},$ it follows that it is also hereditary (cf. [29, Lemma 6.17]; or Subsection 1.2). ∎
@@ 6.3. A sufficient and necessary condition for $\mathcal{PGF}=\mathcal{GP}$
Let $R$ be a ring. Then the following are equivalent::::
@@ 7.1. Application 4: Exact model structures on $\mathcal{GF}_{n}$
The following result is due to R. El Maaouy [22]. It is again a theorem of Šaroch - Št́ovíček type; and when $n=0$ it is given in [28, Theorem 4.11].
@@
([22, Theorem 3.4])  Let $R$ be a ring, $M$ an $R$-module, and $n$ a non-negative integer. Then $\mathcal{PGF}^{\perp}\cap\mathcal{GF}_{n}=\mathcal{F}_{n};$ and the following are equivalent::::
@@
When $n=0$, the following result is [28, Corollary 4.12].
Corollary 7.2.
@@
where $0\leq m\leq n$ (see Lemma 2.2, Lemma 7.3, Corollary 2.6, and Corollary 7.2). Applying Theorem 5.1 to $\mathcal{B}=\mathcal{GF}_{n}$, one gets the assertions (1) - (4) in the following corollary.